OpenStudy (anonymous):

Gail paid $118.65 for a varsity jacket. The price includes a 5% sales tax. How much was the jacket before the tax? $112.72 $113.00 $117.00 $124.65

OpenStudy (anonymous):

118.65 - 118.65(.05) = 118.65 - 5.93 = 112.72